A 1400 kg car starts from rest and accelerates with constant acceleration, traveling 288 m in 13.1 s , what is the force of the road on the car during this acceleration?
1
Expert's answer
2018-10-23T11:30:09-0400

The Newton’s second law states

ma=F

The acceleration of the car

a=2s/t^2

So, the force of the road on the car

F=2ms/t^2 =(2×1400×288)/〖13.1〗^2 =4699 N